# Assumption-lean and Data-adaptive Post-Prediction Inference

### 23 Nov 2023 v1

@@ 1 Introduction
The rest of this paper is organized as follows. In Section 2, we formulate our problem and illustrate our method with mean estimation to build an intuition. Next in Section 3, we introduce a general protocol for applying POP-Inf to estimands defined through estimating equations. Further, we establish the asymptotic properties of the POP-Inf estimator and describe the estimation procedure. We build the connection with the semiparametric efficiency theory in Section 4. Through extensive simulations (Section 5) and real data analysis (Section 6), we validate our theoretical claims and demonstrate the practical utility and effectiveness of the POP-Inf estimator. This paper concludes with a discussion in Section 7.
@@ 2.1 Example: estimating the outcome mean
We provide the intuition of our approach using mean estimation. Consider $U(y,{\bf x};{\bm{\theta}})=y-\theta$. The resulting estimand is $\theta=\mathbb{E}(Y)$, the outcome mean. For illustration propose, we consider ${\bf z}={\bf x}$. Then the classic approach calculates an estimator from the sample average of observed outcomes,
@@
$\widehat{\theta}_{\textnormal{PopInf}}(\omega)=\dfrac{1}{n}\sum_{i=1}^{n}y_{i}+\omega\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}\widehat{f}({\bf z}_{i})\right]$ (3)
@@
$\underbrace{\frac{\textnormal{Var}_{l}(Y)}{n}+\underbrace{\left(\omega^{2}\frac{\textnormal{Var}_{l}[\widehat{f}]}{n}-2\omega\frac{\textnormal{Cov}_{l}[Y,\widehat{f}]}{n}+\omega^{2}\frac{\textnormal{Var}_{u}[\widehat{f}]}{N}\right)}_{\text{Additional terms}}}_{\text{{POP-Inf}\leavevmode\nobreak\ interval}},$ (6)
@@
which achieves its minimum at $\omega_{0}=\textnormal{Cov}_{l}[Y,\widehat{f}({\bf z})]/\{\textnormal{Var}_{l}[\widehat{f}]+n\textnormal{Var}_{l}[\widehat{f}]/N\}$ with minimum value
@@ 3 POP-Inf
$\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+\widehat{\bm{\omega}}\odot\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})\right\}={\bf 0},$ (11)
where $\widehat{\bm{\omega}}$ is an estimate of $\bm{\omega}$ aiming at minimizing the variance. We shall elaborate an optimal value in Section 3.1 and an estimation procedure in Section 3.2. For any weighting vector $\bm{\omega}$, the resulting estimator is denoted as $\widehat{\bm{\theta}}_{\textnormal{PopInf}}(\bm{\omega})$.
@@ 3.1 Asymptotic properties
We study the asymptotic properties of the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{PopInf}}(\bm{\omega})$ with any weight $\bm{\omega}$. To establish its consistency and asymptotic normality, we introduce the following regularity conditions. Condition (C1) is a regularity assumption on the parameter of interest. Conditions (C2)-(C3) guarantee consistency and are weaker than assumptions typically used to guarantee uniform consistency (Newey and McFadden,, 1994). Condition (C4) is needed for asymptotic normality. All conditions are reasonable and standard (Van der Vaart,, 2000).
@@
$\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{PopInf}}({\bm{\omega}})-{\bm{\theta}}_{0})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bm{\Sigma}}(\bm{\omega})),$ (12)
where ${\bm{\Sigma}}(\bm{\omega})={\bf A}^{-1}{\bf V}(\bm{\omega}){\bf A}$, ${\bf A}=\mathbb{E}[\nabla_{{\bm{\theta}}}\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0})]$, ${\bf V}(\bm{\omega})={\bf M}_{1}+\bm{\omega}\bm{\omega}^{\rm T}\odot({\bf M}_{2}+\rho{\bf M}_{3})-2{\bf I}_{q}\bm{\omega}^{\rm T}\odot{\bf M}_{4}$, ${\bf M}_{1}=\textnormal{Var}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{2}=\textnormal{Var}_{l}[\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{3}=\textnormal{Var}_{u}[\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{4}=\textnormal{Cov}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0}),\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$.
@@
${\cal C}_{\alpha,j}=\left\{\theta:\theta\in\left(\widehat{\theta}_{\textnormal{PopInf}}({\bm{\omega}})_{j}-z_{\alpha/2}\{\widehat{\bm{\Sigma}}_{j}{(\bm{\omega})}/n\}^{-1/2},\widehat{\theta}_{\textnormal{PopInf}}({\bm{\omega}})_{j}+z_{\alpha/2}\{\widehat{\bm{\Sigma}}_{j}{(\bm{\omega})}/n\}^{-1/2}\right)\right\},$ (17)
@@ 3.2 Estimation procedure
4:     Update $\widehat{\bm{\omega}}^{k}$ with each element $\widehat{\omega}_{j}^{k}=\mbox{argmin}_{\omega_{j}\in\mathbb{R}}\widehat{\bm{\Sigma}}_{j}(\widehat{\bm{\omega}}^{k-1})\text{ for }j=1,\cdots,q,$.
@@
7:POP-Inf weight $\widehat{\bm{\omega}}$ and the estimator $\widehat{\bm{\theta}}_{\textnormal{PopInf}}(\widehat{\bm{\omega}})$.
@@ 3.3 Examples: linear and logistic regression
$\widehat{\theta}_{\textnormal{PopInf}}({\widehat{\bm{\omega}}})$ =$=\left\{\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}{\bf x}_{i}^{\rm T}+\widehat{\bm{\omega}}\odot\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}{{\bf x}}_{i}{\bf x}_{i}^{\rm T}-\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}{\bf x}_{i}^{\rm T}\right]\right\}^{-1}\cdot\left\{\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}y_{i}+\widehat{\bm{\omega}}\odot\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}{\bf x}_{i}{\widehat{f}}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}\widehat{f}({\bf z}_{i})\right]\right\},$ (21)
@@
For logistic regression, $\bm{\psi}(y,{\bf x};{\bm{\theta}})=-{\bf x}y+{\bf x}\gamma_{{\bm{\theta}}}({\bf x})$, where $\gamma_{{\bm{\theta}}}({\bf x})=1/\left\{1+\exp\left(-{\bf x}^{\rm T}{\bm{\theta}}\right)\right\}$. Using estimation equation (11), the POP-Inf estimator $\widehat{\bm{\theta}}_{\textnormal{PopInf}}({\widehat{\bm{\omega}}})$ isthe solution to the equation
@@ 4 Relationship with Efficient Influence Function
Guided by Theorem 1, an EIF-based estimator would be the solution to the empirical equation,
@@
$\dfrac{1}{n}\sum_{i=1}^{n}\bm{A}^{-1}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+(1-\pi){\bf A}^{-1}\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})\right\}={\bf 0}.$ (26)
@@
Similar to the equation (11) that the POP-Inf estimator solves, the $*{}^{*}$-based estimator (26) solves a equation with two parts: one term that can produce a consistent estimator with labeled data only, and the other term utilizes predictions $\widehat{f}$ that is always mean zero. If one multiples both sides of (11) with matrix ${\bf A}^{-1}$, then (26) corresponds to that equation weight equals to a vector of $1-\pi$. In estimating the outcome mean (Section 2.1) where ${\bf A}$ is essentially one, the $*{}^{*}$-based estimator is the POP-Inf estimator with weight being $1-\pi=N/(n+N)$. The difference between (11) and (26) clarifies one of the main distinctions between the proposed POP-Inf estimator and an EIF guided estimator. The former seeks for variance reduction in a data-adaptive way, while the latter attains minimum variance at correct specifications. In real-world situations, the proposed POP-Inf estimator can be much more efficient in that it utilizes information in a data-adaptive fashion. Numeric comparisons are provided in Section 5.

### 04 Feb 2024 v2

@@ 1 Introduction
The rest of this paper is organized as follows. In Section 2, we formulate our problem and illustrate our method with mean estimation to build an intuition. Next in Section 3, we introduce a general protocol for applying POP-Inf to estimands defined through estimating equations. Further, we establish the asymptotic properties of the POP-Inf estimator and describe the estimation procedure. We build the connection with the semiparametric efficiency theory in Section 4 and describe extensions in Section 5. Through extensive simulations (Section 6) and real data analysis (Section 7), we validate our theoretical claims and demonstrate the practical utility and effectiveness of the POP-Inf estimator. This paper concludes with a discussion in Section 8.
@@ 2.1 Example: estimating the outcome mean
We provide the intuition of our approach using mean estimation. Consider $\psi(y,{\bf x};{\bm{\theta}})=y-\theta$. The resulting estimand is $\theta=\mathbb{E}(Y)$, the outcome mean. For illustration propose, we consider ${\bf z}={\bf x}$. Then the classic approach calculates an estimator from the sample average of observed outcomes,
@@
$\widehat{\theta}_{\textnormal{POP-Inf}}(\omega)=\dfrac{1}{n}\sum_{i=1}^{n}y_{i}+\omega\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}\widehat{f}({\bf z}_{i})\right]$ (3)
@@
$\underbrace{\frac{\textnormal{Var}_{l}(Y)}{n}+\underbrace{\left(\omega^{2}\frac{\textnormal{Var}_{l}[\widehat{f}]}{n}-2\omega\frac{\textnormal{Cov}_{l}[Y,\widehat{f}]}{n}+\omega^{2}\frac{\textnormal{Var}_{u}[\widehat{f}]}{N}\right)}_{\text{Additional terms}}}_{\text{Variance of {POP-Inf}\leavevmode\nobreak\ estimator}},$ (6)
@@
which achieves its minimum at $\omega_{0}=\textnormal{Cov}_{l}[Y,\widehat{f}({\bf z})]/\{\textnormal{Var}_{l}[\widehat{f}]+n\textnormal{Var}_{u}[\widehat{f}]/N\}$ with minimum value
@@ 3 POP-Inf
$\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+\bm{\omega}\odot\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})\right\}={\bf 0},$ (11)
where $\bm{\omega}$ can be replaced with an estimate $\widehat{\bm{\omega}}$ aiming at minimizing the variance. We shall elaborate an optimal value in Section 3.1 and an estimation procedure in Section 3.2. For any weighting vector $\bm{\omega}$, the resulting estimator is denoted as $\widehat{\bm{\theta}}_{\textnormal{POPInf}}(\bm{\omega})$.
@@ 3.1 Asymptotic properties
We study the asymptotic properties of the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{POPInf}}(\bm{\omega})$ with any weight $\bm{\omega}$. To establish its consistency and asymptotic normality, we introduce the following regularity conditions. Condition (C1) is a regularity assumption on the parameter of interest. Conditions (C2)-(C3) guarantee consistency. Condition (C4) is needed for asymptotic normality. All conditions are reasonable and standard (Van der Vaart,, 2000).
@@
$\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}({\bm{\omega}})-{\bm{\theta}}_{0})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bm{\Sigma}}(\bm{\omega})),$ (12)
where ${\bm{\Sigma}}(\bm{\omega})={\bf A}^{-1}{\bf V}(\bm{\omega}){\bf A}^{-1}$, ${\bf A}=\mathbb{E}[\partial\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0})/\partial{\bm{\theta}}]$, ${\bf V}(\bm{\omega})={\bf M}_{1}+\bm{\omega}\bm{\omega}^{\rm T}\odot({\bf M}_{2}+\rho{\bf M}_{3})-2{\bf I}_{q}\bm{\omega}^{\rm T}\odot{\bf M}_{4}$, ${\bf M}_{1}=\textnormal{Var}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{2}=\textnormal{Var}_{l}[\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{3}=\textnormal{Var}_{u}[\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{4}=\textnormal{Cov}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0}),\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$. Proof is contained in Appendix B.1.
@@
${\cal C}_{\alpha,j}=\left\{\theta:\theta\in\left(\widehat{\theta}_{\textnormal{POP-Inf}}({\bm{\omega}})_{j}-z_{\alpha/2}\{\widehat{\bm{\Sigma}}_{j}{(\bm{\omega})}/n\}^{-1/2},\widehat{\theta}_{\textnormal{POP-Inf}}({\bm{\omega}})_{j}+z_{\alpha/2}\{\widehat{\bm{\Sigma}}_{j}{(\bm{\omega})}/n\}^{-1/2}\right)\right\},$ (17)
@@ 3.2 Estimation procedure
4:     Update $\widehat{\bm{\omega}}^{k}$ with each element $\widehat{\omega}_{j}^{k}=\mbox{argmin}_{\omega_{j}\in\mathbb{R}}\widehat{\bm{\Sigma}}_{j}(\widehat{\bm{\omega}}^{k-1})\text{ for }j=1,\cdots,q$.
@@
7:POP-Inf weight $\widehat{\bm{\omega}}$ and the estimator $\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}(\widehat{\bm{\omega}})$.
@@ 3.3 Examples: linear and logistic regression
$\widehat{\theta}_{\textnormal{POP-Inf}}({\widehat{\bm{\omega}}})$ =$=\left\{\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}{\bf x}_{i}^{\rm T}+\widehat{\bm{\omega}}\odot\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}{{\bf x}}_{i}{\bf x}_{i}^{\rm T}-\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}{\bf x}_{i}^{\rm T}\right]\right\}^{-1}\cdot\left\{\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}y_{i}+\widehat{\bm{\omega}}\odot\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}{\bf x}_{i}{\widehat{f}}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}\widehat{f}({\bf z}_{i})\right]\right\},$ (21)
@@
For logistic regression, $\bm{\psi}(y,{\bf x};{\bm{\theta}})=-{\bf x}y+{\bf x}\gamma_{{\bm{\theta}}}({\bf x})$, where $\gamma_{{\bm{\theta}}}({\bf x})=1/\left\{1+\exp\left(-{\bf x}^{\rm T}{\bm{\theta}}\right)\right\}$. Using estimation equation (11), the POP-Inf estimator $\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}({\widehat{\bm{\omega}}})$ isthe solution to the equation
@@
3.4 Interpretation of the weighting vector
Next, we provide interpretations of the weighting vector. We build the intuitions using two simple examples, mean estimation and simple linear regression with one-dimensional covariates.
@@ 4 Relationship with Efficient Influence Function
Proof is contained in Appendix B.2.
Guided by Proposition 1, an EIF-based estimator would be the solution to the empirical equation,
@@
$\dfrac{1}{n}\sum_{i=1}^{n}{\bf A}^{-1}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+(1-\pi){\bf A}^{-1}\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})\right\}={\bf 0}.$ (26)
@@
Similar to the equation (11) that the POP-Inf estimator solves, the $*{}^{*}$-based estimator (26) solves a equation with two parts: one term that can produce a consistent estimator with labeled data only, and the other term utilizes predictions $\widehat{f}$ that is always mean zero. If one multiples both sides of (11) with matrix ${\bf A}^{-1}$, then (26) corresponds to that equation weight equals to a vector of $1-\pi$. In estimating the outcome mean (Section 2.1) where ${\bf A}$ is essentially the scalar one, the $*{}^{*}$-based estimator is the POP-Inf estimator with weight being $1-\pi=N/(n+N)$. The difference between (11) and (26) clarifies one of the main distinctions between the proposed POP-Inf estimator and an EIF guided estimator. The former seeks for variance reduction in a data-adaptive way, while the latter attains minimum variance at correct specifications. In real-world situations, the proposed POP-Inf estimator can be much more efficient in that it utilizes information in a data-adaptive fashion. Numeric comparisons are provided in Section 6.
5 Extensions
Here, we make extensions of POP-Inf to handle more general situations. In particular, Section 5.1 describes an inference procedure when the the outcome is fully labeled but covariate ${\bf x}$ is partly labeled and Section 5.2 considers the case where both $y$ and ${\bf x}$ are only partly labeled.
5.1 ML-predicted covariates
We adjusted the previously described setup by modifying it to use ML to predict the covariate ${\bf X}$ instead of the outcome $Y$. Our data can be divided into two parts: ${\cal L}\cup{\cal U}$. The ${\cal L}$ part includes data points $(y_{i},{\bf x}_{i},{\bf z}_{i})$ for $i=1,\cdots,n$ and the ${\cal U}$ part includes data points $(y_{i},{\bf z}_{i})$ for $i=n+1,\cdots,N+n$. In addition, we use an external prediction algorithm with the notation $\widehat{q}(\cdot)$. This algorithm is applied to ${\bf z}$ to generate the predicted value $\widehat{q}({\bf z})$ of the covariate ${\bf X}$. However, it is important to note that $\widehat{q}$ may produce inaccurate or biased predictions of ${\bf X}$. The estimand is also defined by the estimating equation (1).
Similar to our proposal to handle the ML-predicted outcome, we propose our POP-Inf estimator with ML-predicted covariates that solves the equation,
$\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+\widehat{\bm{\omega}}\odot\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},\widehat{q}({\bf z}_{i});{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(y_{i},\widehat{q}({\bf z}_{i});{\bm{\theta}})\right\}={\bf 0},$ (27)
where $\widehat{\bm{\omega}}$ is also to minimize the variance of the estimator. Such estimation equation is also the estimation equation for the classic estimator with an augmented mean zero term that is indexed by a vector $\bm{\omega}$. Therefore, the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{POPInf}}(\bm{\omega})$ is also consistent and asymptotic normally distributed, under mild regularity conditions. Corollary 3 states the asymptotic distribution that enables interval estimation with its proof follows from the proof of Theorem 1.
Corollary 3.
$\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}({\bm{\omega}})-{\bm{\theta}}_{0})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bm{\Sigma}}^{\prime}(\bm{\omega})).$ (28)
where ${\bm{\Sigma}}^{\prime}(\bm{\omega})={\bf A}^{-1}{\bf V}^{\prime}(\bm{\omega}){\bf A}^{-1}$, ${\bf V}^{\prime}(\bm{\omega})={\bf M}_{1}+\bm{\omega}\bm{\omega}^{\rm T}\odot({\bf M}_{2}^{\prime}+\rho{\bf M}_{3}^{\prime})-2{\bf I}_{q}\bm{\omega}^{\rm T}\odot{\bf M}_{4}^{\prime}$, ${\bf M}_{2}^{\prime}=\textnormal{Var}_{l}[\bm{\psi}(Y,\widehat{q}({\bf X});{\bm{\theta}}_{0})]$, ${\bf M}_{3}=\textnormal{Var}_{u}[\bm{\psi}(Y,\widehat{q}({\bf X});{\bm{\theta}}_{0})]$, ${\bf M}_{4}^{\prime}=\textnormal{Cov}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0}),\bm{\psi}(Y,\widehat{q}({\bf X});{\bm{\theta}}_{0})]$.
5.2 ML-predicted outcome and covariates
Next, we explore a situation where the outcome and covariates are not directly observed in the unlabeled data. Instead, we employ machine learning techniques to predict both. Additionally, we possess a much smaller data with both outcome and covariates measured. This full data can be categorized into two segments: ${\cal L}\cup{\cal U}$. The first segment, ${\cal L}$, includes the data points $(y_{i},{\bf x}_{i},{\bf z}_{i})$ for $i=1,\cdots,n$, while the second segment, ${\cal U}$, comprises the instances of ${\bf z}_{i}$ for $i=n+1,\cdots,N+n$, and $N>>n$. The variable ${\bf z}$ is to predict both the outcome and the covariates. The ML-predicted outcome and covariates are represented by $\widehat{f}({\bf z})$ and $\widehat{q}({\bf z})$, respectively, where $\widehat{f}$ and $\widehat{q}$ are independently obtained. The POP-Inf estimator for this situation is to solve the equation,
$\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+\widehat{\bm{\omega}}\odot\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),\widehat{q}({\bf z}_{i});{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),\widehat{q}({\bf z}_{i});{\bm{\theta}})\right\}={\bf 0},$ (29)
where $\widehat{\bm{\omega}}$ can be obtained by minimizing the variance of the estimator. Similarly, we have the consistency and asymptotic normality for the $\widehat{\bm{\theta}}_{\textnormal{POPInf}}(\bm{\omega})$ in Corollary 4 whose proof follows from the proof of Theorem 1.
Corollary 4.
$\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}({\bm{\omega}})-{\bm{\theta}}_{0})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bm{\Sigma}}^{\prime\prime}(\bm{\omega})),$ (30)
where ${\bm{\Sigma}}^{\prime\prime}(\bm{\omega})={\bf A}^{-1}{\bf V}^{\prime\prime}(\bm{\omega}){\bf A}^{-1}$, ${\bf V}^{\prime\prime}(\bm{\omega})={\bf M}_{1}+\bm{\omega}\bm{\omega}^{\rm T}\odot({\bf M}^{\prime\prime}_{2}+\rho{\bf M}^{\prime\prime}_{3})-2{\bf I}_{q}\bm{\omega}^{\rm T}\odot{\bf M}^{\prime\prime}_{4}$, ${\bf M}_{2}^{\prime\prime}=\textnormal{Var}_{l}[\bm{\psi}(\widehat{f}({\bf Z}),\widehat{q}({\bf X});{\bm{\theta}}_{0})]$, ${\bf M}_{3}^{\prime\prime}=\textnormal{Var}_{u}[\bm{\psi}(\widehat{f}({\bf Z}),\widehat{q}({\bf X});{\bm{\theta}}_{0})]$, ${\bf M}_{4}^{\prime\prime}=\textnormal{Cov}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0}),\bm{\psi}(\widehat{f}({\bf Z}),\widehat{q}({\bf X});{\bm{\theta}}_{0})]$.